Jolene invests her savings in two bank accounts, one paying 5 percent and the other paying 12 percent simple interest per year. She puts twice as much in the lower-yielding account because it is less risky. Her annual interest is 4840 dollars. How much did she invest at each rate?

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52781)

.

Let x = the 12% investment amount.

then the 5% investment is 2x.



The total annual interest equation is

    0.5*(2x) + 0.12x = 4840  dollars.



You do the rest.



Answer by greenestamps(13200)



An amount 2x invested at 5% is the same as an amount x invested at 10%.

So the problem is equivalent to investing equal amounts at 10% and 12%.

That means the average interest rate is 11%.

$4840 interest at 11% means the total investment was $4840/0.11 = $44,000.

So the investment was equivalent to $22,000 at 10% and $22,000 at 12%. But since the interest rates were actually 5% and 12%, the amounts invested at each rate were

ANSWERS: $44,000 at 5% and $22,000 at 12%.

CHECK: .05(44000)+.12(22000) = 2200+2640 = 4840